In a constant flow experiment, a liquid flows at the rate of 0.15kg/min through a tube and is heated dissipating 25.2W. the inflow and outflow temperatures are 15.2 degree Celsius and 17.4 degree Celsius respectively. When the rate of flow is increased to 0.2318kg/min and the rate of heating to 37.8W, the inflow and outflow temperatures remain unaltered. Find the Specific Heat Capacity of the liquid
1
Expert's answer
2021-02-25T18:05:53-0500

"mass rate_1 = 0.15kg/min = 0.0025 kg/sec\\

P = 25.2W\\

∆T = 17.4 - 15.2 = 2.2°C"


"mass\ rate_2 = 0.2318kg/min = 0.00386 kg/sec\\

P = 37.8W
\\
∆T = 17.4 - 15.2 = 2.2°C"


from,

"Pt = mc∆\theta\\
P = \dfrac{mc∆\theta}{t} = \dfrac{m}{t} c∆\theta"


however, Total Power = "\dfrac{m}{t} c∆\theta" + heat dissipated


"25.2 = 0.0025 × c × 2.2 + H\\

25.2 = 0.0055c + h ---(i)"


"37.8 = 0.00386 × c × 2.2 + H\\

37.8 = 0.00849c + h ---(ii)"


subtracting equation (ii) from (i)

"12.6 = 0.00299c\\

c = 4214.05 J/kg.K"